# When Tokens Talk Too Much: A Survey of Multimodal Long-Context Token Compression across Images, Videos, and Audios

### v4

@@ Abstract
Index Terms:
Multimodal Large Language Models, Token Compression, Long Context, Efficient AI
I Introduction
Multimodal large language models (MLLMs) [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11] have demonstrated exceptional performance on complex tasks, including visual question answering (VQA), automatic speech recognition (ASR) and multimodal content generation, by extending the architectural principles of large language models (LLMs) [12, 13, 14, 15]. These powerful models derive their strength from processing very long and diverse contexts, such as high-resolution images, extended video sequences, and long audio input, using transformer architectures.
Figure 1: Up: Image, video, and audio data types can scale in their representation dimensions, leading to a corresponding increase in the number of tokens. Down: Top-performing MLLMs cannot address real-world demands, as the number of tokens for multimodal input, especially video, vastly exceeds that of text, and most visual tokens are redundant. Therefore, token compression is crucial to address this limitation.
Achieving this capability, however, faces a significant challenge: the quadratic complexity of the self-attention mechanism. As the number of tokens increases, this complexity leads to substantial computational and memory demands. This problem is particularly pronounced in MLLMs, where the tokenization of visual and audio data can generate sequences of orders of magnitude longer than text [16, 17, 18].
@@ I Introduction
To address the challenges posed by the long context, token compression has emerged as a critical research focus for enhancing the inference efficiency and practical deployment of MLLMs. This approach is highly effective because multimodal inputs, like those processed by vision transformers (ViT), contain significant redundancy [19, 20, 21, 22, 23, 24, 25]. Extensive research, for example, demonstrates that more than $50\%$ of tokens in a typical MLLM sequence receive minimal attention during inference [26, 27, 17, 16, 28, 29]. While some advanced techniques integrate compression directly into a model’s architecture or training framework [30, 31, 32, 33, 2, 34, 35, 36, 37, 38, 39, 40], a major advantage of token compression is its ability to be applied as a post-optimization technique without requiring expensive retraining. These methods typically operate by first establishing a specialized metric to evaluate token importance, then performing a corresponding pruning or compression. By significantly accelerating inference and reducing memory consumption, these techniques enable the practical deployment of MLLMs in real-world applications [41, 38, 39, 42, 43].
@@
Recent extensive research demonstrates that token compression substantially enhances inference efficiency, driving the continuous development of diverse compression strategies and sophisticated methodologies [44, 45, 28, 27, 46, 18, 29, 47, 48, 49, 26, 50, 51, 52, 25]. However, the inherent heterogeneity of multimodal data means that redundancy manifests differently across modalities. Unlike textual prompts, where redundancy is primarily in syntactic or semantic, visual and auditory data exhibit unique structural properties. For instance, high-resolution images contain strong local correlations, while video streams feature extensive spatiotemporal redundancy across frames, and audio signals often contain extended segments of silence or stationary noise. Consequently, most existing methods focus on compressing one or two specific modalities.
@@
Significant strides have been made in compressing tokens in text LLMs. For instance, [53] has thoroughly explored prompt compression for text LLMs, highlighting advancements in this domain. In MLLMs, position paper [54] has begun to broaden our understanding, emphasizing that token compression offers benefits beyond mere efficiency. Furthermore, some researchers argue that the focus of research for efficient AI is shifting from model-centric compression to data-centric compression [55]. However, there has not yet been a systematic classification of token compression methods specifically for MLLMs, leaving an opportunity for a comprehensive survey in this area.
@@
The remaining sections of the article are organized as follows: we will first discuss the architecture of MLLMs in the background section (Section II-A), followed by an examination of how token compression has been utilized in prior methods for large language models (LLMs, Section II-B) and vision transformers (ViTs, Section II-C). Subsequent sections will be dedicated to token compression methods for specific modalities: Section III for image LLMs, Section IV for video LLMs, and Section V for audio LLMs. Following this, Section VI will provide insights into token compression research. Finally, Section VII will introduce the broad application space of token compression, followed by the concluding Section VIII.
@@
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-black, rounded corners, align=left, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=4pt, line width=1.1pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=10.5em,font=,, where level=2text width=11.5em,font=,, where level=3text width=12em,font=,, where level=4text width=50em,font=,, [Multimodal Token Compression, ver, [    Image (§III)    ,ver [ Transformation-based
    (§III-A)    [ InternVL1.5 [30], NVLM [31], Qwen2-VL [32], Qwen2.5-VL [33], LaCo [56], LLaVA-OneVision [2],
LLaVA-Video [34], Seed1.5-VL [57], M3 [35], DeCo [36], SlowFast-LLaVA [58], PruMerge+ [29],
C-Abstractor [37], MobileVLM [38], MobileVLM V2 [39], LLaMA-VID [40], Dynamic-VLM [59] , leaf, text width=44em] ] [   Similarity-based
    (§III-B)    [ ToMe [60], FOLDER [61], DivPrune [28], AuroraCap [45], TopV [49], Skip-Vision [62], DyMU [63],
Dynamic-VLM [59], STTM [64], VisionZip [18], VisPruner [65], PuMer [48], iLLaVA [66],
G-Prune [67], BTP [68] , leaf, text width=44em] ] [   Attention-based
    (§III-C)    [ Attention in Encoder: PruMerge+ [29], VisionZip [18], VisPruner [65], GlobalCom2 [69],
MustDrop [52], VScan [47], HiRED [70], FiCoCo [71]
Attention in Decoder: FastV [26], PyramidDrop [72], VTW [73], FitPrune [74], ST3 [75],
ATP-LLaVA [76], CoreMatching [77], ZipVL [46], HoliTom [16], TokenCarve [78], SparseVLM [51],
MustDrop [52], VScan [47], FrameFusion [79], iLLaVA [66], G-Search [80], FiCoCo [71], BTP [68] , leaf, text width=44em] ] [   Query-based
    (§III-D)    [ Token Distillation: InstructBLIP [1], BLIP-2 [81], mPLUG-Owl [82], Minigpt-4 [83], Flamingo [84],
Qwen-VL [4], LLaMA-VID [40], LLaVA-Mini [85], VoCo-LLaMA [86], Victor [87]
Cross-Modal Selection: SparseVLM [51], AdaFV [88], TRIM [89] , leaf, text width=44em] ] ] [    Video (§IV)    , ver [ Transformation-based
    (§IV-A)    [ PLLaVA [3], Video-ChatGPT [90], LongVLM [91], VideoLLaMA 2 [10], SlowFast-LLaVA-1.5 [92],
VideoChat-Flash [93], MobileVLM [38] , leaf3, text width=44em] ] [   Similarity-based
    (§IV-B)    [ Chat-UniVi [94], PruneVid [27], FastVID [44], HoliTom [16], DyCoke [17], FrameFusion [79],
LongVU [95], VidCom2 [96], DynTok [97], MovieChat [98], VideoChat-Flash [93], Video-XL [99],
TimeChat-Online [100], DTC [101], LLaVA-Scissor [102], Aim [103] , leaf3, text width=44em] ] [   Attention-based
    (§IV-C)    [ Attention in Encoder: PruMerge+ [29], VisionZip [18], VisPruner [65], MustDrop [52], FiCoCo [71]
Attention in Decoder: FastV [26], PyramidDrop [72], VTW [73], CoreMatching [77], ZipVL [46],
SparseVLM [51], FastVID [44], MustDrop [52], HoliTom [16], VScan [47], FrameFusion [79],
iLLaVA [66], G-Search [80], FiCoCo [71], Aim [103] , leaf3, text width=44em] ] [   Query-based
    (§IV-D)    [ Token Distillation: Token Turing Machines [104], BLIP-3-Video [105], Long-VMNet [106],
STORM [107], LinVT [108] LLaMA-VID [40], LLaVA-Mini [85], VideoChat-Flash [93]
Cross-Modal Selection: LongVU [95] , leaf3, text width=44em] ] ] [    Audio (§V)    , ver [ Transformation-based
    (§V-A)    [ HTS-AT [109], SLAM-ASR [110], LLaMA-Omni [111], SpeechVerse [112], Qwen2-audio [113],
Qwen2.5-Omni [5], Baichuan-Audio [114], Llama-AVSR [115], Llama-MTSK [116], OSUM [117],
LUCY [118] , leaf5, text width=44em] ] [   Similarity-based
    (§V-B)    [ A-ToMe [119] , leaf5, text width=44em] ] [   Attention-based
    (§V-C)    [ Attention in Encoder: Top-K [120]
Attention in Decoder: SpeechPrune [121] , leaf5, text width=44em] ] [   Query-based
    (§V-D)    [ Token Distillation: Video-LLaMA [7], SALMONN [122], video-SALMONN [123], Typhoon 2 [124],
MMCE-Qformer [125], MMS-LLaMA [126]
Cross-Modal Selection: SpeechPrune [121] , leaf5, text width=44em] ] ] ]
@@
II Background
II-A Multimodal Architecture
@@
  • Vision encoders: CLIP [127], SigLIP [128], DINO [129, 130], and ViT [33];
  • Audio encoders: Whisper [131] and Audio-CLIP [132].
@@
The pre-trained LLM [12, 13, 14] forms the core of the framework, with its large-scale parameters providing emergent capabilities such as zero-shot generalization and in-context learning. The LLM receives a composite input sequence formed by concatenating the projected multimodal embeddings $H_{v}$ and $H_{a}$, as well as the textual prompt embeddings $H_{q}$. The textual prompt $X_{q}$ is first converted into embeddings $H_{q}$ by an integrated tokenizer.
@@
The high dimensionality of multimodal data poses a computational challenge. As shown in Figure 2, the token sequence processed comprises a mix of system prompt, multimodal context, and textual instruction. In most reasoning tasks, multimodal tokens constitute over $80\%$ of the total sequence length [26], thereby forming the primary computational bottleneck. This bottleneck obstacle to scaling MLLMs and achieving efficient inference. Consequently, a key strategy to optimize computational efficiency involves employing specialized projector architectures. These projectors are designed to reduce the number of multimodal tokens while preserving their semantic fidelity, thus mitigating the computational burden.
@@
While MLLM architecture presents unique challenges, token compression has been explored for both encoders and LLMs independently. Therefore, the subsequent sections will first dive into techniques relevant to these individual components, paving the way for more efficient multimodal models. Specifically, Section II-B will focus on token compression methods for large language models (LLMs), and Section II-C will explore techniques for vision transformers (ViTs).
II-B Large Language Model Token Compression
@@
The backbone of modern MLLMs is often built upon and fine-tuned from powerful text-based LLMs. As a foundational component, a solid understanding of token compression techniques developed for text LLMs is crucial, as they offer an accurate and lightweight solution for handling real-world long-context scenarios, such as understanding an entire book or a code repository. Within the domain of large language models, these methods are frequently termed prompt compression [53].
@@
AutoCompressor [133] condenses context into summary vectors as soft prompts. Extensible Tokenization [134] employs intermediate modules to compress embeddings, while SentenceVAE [135] represents sentences with single tokens. Selective Context [136] employs self-information metrics to eliminate low-information tokens. LLMLingua [137, 138, 139] series utilizes hierarchical token pruning with instruction tuning and further introduces LongLLMLingua [138] to mitigate position decay through semantic density ranking. In parallel, query-guided methods like QUITO [140] and QUITO-X [141] leverage attention scores or information bottleneck theory for relevance-based filtering. AdaComp [142] implements adaptive extraction governed by query complexity predictors. Concept Distillation [143] employs Abstract Meaning Representation (AMR) graphs to distill key concepts, whereas xRAG [144] collapses documents into single-token representations. ICAE [145] encodes context into discrete memory slots. Recursive frameworks including RCC [146] and XL3M [147] generate piecewise summaries through relevant fusion. SoftPromptComp [148] fuses natural language prompts with dynamic embeddings, while PromptIntern [149] internalizes task instructions into model parameters via phased training.
@@
While these text-centric token compression techniques have demonstrated notable efficacy, their direct application to MLLMs faces fundamental challenges. The inherent heterogeneity of multimodal data introduces distinct redundancy patterns that are not present in unimodal text. These include, but are not limited to, spatial correlations in high-resolution images, spatiotemporal continuity in video sequences, and spectral-temporal locality in audio streams. These specialized redundancies necessitate the development of dedicated compression strategies. Consequently, this survey systematically reviews emerging token compression methodologies designed specifically for MLLMs that effectively reduce token redundancy while preserving task performance.
II-C Vision Transformer Token Compression
@@
Visual token compression, originally pioneered in vision transformers (ViTs) [24, 150, 151, 152, 153, 154, 155, 156], offers insights for addressing analogous challenges in MLLMs.
@@
Spatial redundancy manifests in ViTs through adjacent image patches, where not all tokens contribute equally to classification outcomes, compounded by semantic imbalance: foreground objects demand disproportionate computational resources compared to homogeneous backgrounds. To mitigate these issues, visual token compression techniques are employed to reduce computational overhead while maintaining model accuracy substantially.
@@
Foundational approaches, including DynamicViT [19] and EViT [20], quantify token relevance through attention scores, dynamically pruning low-saliency tokens. Complementary techniques like ToMe [21] and TokenLearner [22] either merge semantically similar tokens using similarity metrics or generate compact token sets via learned spatial attention mechanisms. DeiT [23] employs lightweight ‘student’ heads to predict categorical labels from compressed token subsets. Furthermore, methods such as MADTP [157] leverage cross-modal alignment to filter tokens.
@@
TABLE I: Four Categories of Methods Based on Intrinsic Mechanisms: Diagram, Summary, and Pros & Cons.

### v5

# A Survey of Token Compression for Efficient Multimodal Large Language Models

@@ Abstract
1 Introduction
Multimodal large language models (MLLMs) (Liu et al., 2023; Li et al., 2025a; Xu et al., 2024; Bai et al., 2023; Xu et al., 2025b; Lin et al., 2024a; Zhang et al., 2023a; Li et al., 2024a; 2023d; Cheng et al., 2024c; Zhang et al., 2025a; Song et al., 2024) have demonstrated exceptional performance on complex tasks, including visual question answering (VQA), automatic speech recognition (ASR) and multimodal content generation, by extending the architectural of large language models (LLMs) (Chiang et al., 2023; Team, 2024; AI@Meta, 2024; Abdin et al., 2024). These powerful models derive their strength from processing long and diverse contexts, such as high-resolution images, extended video sequences, and long audio input, using transformer architectures.
Refer to caption Refer to caption
Figure 1: Left: Image, video, and audio data types can scale in their representation dimensions, leading to a corresponding increase in the number of tokens. Right: Top-performing MLLMs cannot address real-world demands, as the number of tokens for multimodal input, especially video, vastly exceeds that of text, and most visual tokens are redundant. Therefore, token compression is crucial to address this limitation.
Achieving this capability, however, faces a significant challenge: the quadratic complexity of the self-attention mechanism. As the number of tokens increases, this complexity leads to substantial computational and memory demands. This problem is particularly pronounced in MLLMs, where the tokenization of visual and audio data can generate sequences of orders of magnitude longer than text (Shao et al., 2025; Tao et al., 2025a; Yang et al., 2025c; Song et al., 2025c).
@@ 1 Introduction
To address the challenges posed by the long context, token compression has emerged as a critical research focus for enhancing the inference efficiency and practical deployment of MLLMs. This approach is highly effective because multimodal inputs, like those processed by vision transformers (ViT), contain significant redundancy (Rao et al., 2021; Liang et al., 2022; Bolya et al., 2022; Ryoo et al., 2021; Touvron et al., 2021; Vaswani et al., 2017; Yang et al., 2025d). Extensive research, for example, demonstrates that more than $50\%$ of tokens in a typical MLLM sequence receive minimal attention during inference (Chen et al., 2024a; Huang et al., 2025b; Tao et al., 2025a; Shao et al., 2025; Alvar et al., 2025; Shang et al., 2025). While some advanced techniques integrate compression directly into a model’s architecture or training framework (Chen et al., 2024c; Dai et al., 2024; Wang et al., 2024b; Bai et al., 2025; Li et al., 2025a; Zhang et al., 2024d; Cai et al., 2024a; Yao et al., 2024; Cha et al., 2024; Chu et al., 2023; 2024a; Li et al., 2024c), a major advantage of token compression is its ability to be applied as a post-optimization technique without requiring expensive retraining. These methods typically operate by first establishing a specialized metric to evaluate token importance, then performing a corresponding pruning or compression. By significantly accelerating inference and reducing memory consumption, these techniques enable the practical deployment of MLLMs in real-world applications (Lin et al., 2025a; Chu et al., 2023; 2024a; Wei et al., 2025; Ma et al., 2024b).
@@
Recent extensive research demonstrates that token compression substantially enhances inference efficiency, driving the continuous development of diverse compression strategies and sophisticated methodologies (Shen et al., 2025a; Chai et al., 2025; Alvar et al., 2025; Huang et al., 2025b; Yang et al., 2025c; Shang et al., 2025; Zhang et al., 2025b; Cao et al., 2023; Yang et al., 2025a; Chen et al., 2024a; Tao et al., 2025c; Zhang et al., 2024c; Liu et al., 2024c; Yang et al., 2025d; Ma et al., 2025b). However, the inherent heterogeneity of multimodal data means that redundancy differs across modalities. Unlike textual prompts, where redundancy is primarily in syntactic or semantic, visual and auditory data exhibit unique structural properties. For instance, high-resolution images contain strong local correlations, while video streams feature extensive spatiotemporal redundancy across frames, and audio signals often contain extended segments of silence or stationary noise. Consequently, most existing methods focus on compressing one or two specific modalities.
@@
Significant strides have been made in compressing tokens in text LLMs. For instance, (Li et al., 2025c) has thoroughly explored prompt compression for text LLMs, highlighting advancements in this domain. In MLLMs, position paper (Kong et al., 2025) has begun to broaden our understanding, emphasizing that token compression offers benefits beyond mere efficiency. Furthermore, some researchers argue that the focus of research for efficient AI is shifting from model-centric compression to data-centric compression (Liu et al., 2025b). However, there has not yet been a systematic classification of token compression methods specifically for MLLMs, leaving an opportunity for a comprehensive survey in this area.
@@
The remaining sections of the article are organized as follows: we will first discuss the architecture of MLLMs in the background section (Section 2.1), followed by an examination of how token compression has been utilized in prior methods for large language models (LLMs, Section 2.2) and vision transformers (ViTs, Section 2.3). Subsequent sections will be dedicated to token compression methods for specific modalities: Section 3 for image LLMs, Section 4 for video LLMs, and Section 5 for audio LLMs. Following this, Section 6 will provide insights into token compression research. Finally, Section 7 will introduce the broad application space of token compression, followed by the concluding Section 8.
@@
2 Background
2.1 Multimodal Architecture
@@
  • Vision encoders: CLIP (Radford et al., 2021), SigLIP (Zhai et al., 2023), DINO (Caron et al., 2021; Oquab et al., 2023), and ViT (Bai et al., 2025);
  • Audio encoders: Whisper (Radford et al., 2023) and Audio-CLIP (Guzhov et al., 2022).
@@
The pre-trained LLM (Chiang et al., 2023; Team, 2024; AI@Meta, 2024) forms the core of the framework, with its large-scale parameters providing emergent capabilities such as zero-shot generalization and in-context learning. The LLM receives a composite input sequence formed by concatenating the projected multimodal embeddings $H_{v}$ and $H_{a}$, as well as the textual prompt embeddings $H_{q}$. The textual prompt $X_{q}$ is first converted into embeddings $H_{q}$ by an integrated tokenizer.
@@
The high dimensionality of multimodal data poses a computational challenge. As shown in Figure 2, the token sequence processed comprises a mix of system prompt, multimodal context, and textual instruction. In most reasoning tasks, multimodal tokens constitute over $80\%$ of the total sequence length (Chen et al., 2024a), thereby forming the primary computational bottleneck. This bottleneck obstacle to scaling MLLMs and achieving efficient inference. Consequently, a key strategy to optimize computational efficiency involves employing specialized projector architectures. These projectors are designed to reduce the number of multimodal tokens while preserving their semantic fidelity, thus mitigating the computational burden.
@@
While MLLM architecture presents unique challenges, token compression has been explored for both encoders and LLMs independently. Therefore, the subsequent sections will first dive into techniques relevant to these individual components, paving the way for more efficient multimodal models. Specifically, Section 2.2 will focus on token compression methods for large language models (LLMs), and Section 2.3 will explore techniques for vision transformers (ViTs).
2.2 Large Language Model Token Compression
@@
The backbone of modern MLLMs is often built upon and fine-tuned from powerful text-based LLMs. As a foundational component, a solid understanding of token compression techniques developed for text LLMs is crucial, as they offer an accurate and lightweight solution for handling real-world long-context scenarios, such as understanding an entire book or a code repository. Within the domain of large language models, these methods are frequently termed prompt compression (Li et al., 2025c).
@@
AutoCompressor (Chevalier et al., 2023) condenses context into summary vectors as soft prompts. Extensible Tokenization (Shao et al., 2024) employs intermediate modules to compress embeddings, while SentenceVAE (An et al., 2024) represents sentences with single tokens. Selective Context (Li et al., 2023g) employs self-information metrics to eliminate low-information tokens. LLMLingua (Jiang et al., 2023a; b; Pan et al., 2024) series utilizes hierarchical token pruning with instruction tuning and further introduces LongLLMLingua (Jiang et al., 2023b) to mitigate position decay through semantic density ranking.
In parallel, query-guided methods like QUITO (Wang et al., 2024d) and QUITO-X (Wang et al., 2024e) leverage attention scores or information bottleneck theory for relevance-based filtering. AdaComp (Zhang et al., 2024a) implements adaptive extraction governed by query complexity predictors. Concept Distillation (Shi et al., 2024) employs Abstract Meaning Representation (AMR) graphs to distill key concepts, whereas xRAG (Cheng et al., 2024b) collapses documents into single-token representations. ICAE (Ge et al., 2023) encodes context into discrete memory slots. Recursive frameworks including RCC (Huang et al., 2024) and XL3M (Wang et al., 2024c) generate piecewise summaries through relevant fusion. SoftPromptComp (Wang et al., 2024a) fuses natural language prompts with dynamic embeddings, while PromptIntern (Zou et al., 2024) internalizes task instructions into model parameters via phased training.
Targeting inference efficiency, KV cache compression techniques prune redundant memory states to accelerate generation. H2O (Zhang et al., 2023b) and StreamingLLM (Xiao et al., 2024) utilize heavy-hitter policies and attention sinks to maintain generation quality under limited budgets. Furthermore, SnapKV (Li et al., 2024d) and PyramidKV (Cai et al., 2024b) enhance long-context performance by pinpointing key attention clusters or dynamically adjusting cache allocations across layers.
@@
While these text-centric token compression techniques have demonstrated notable efficacy, their direct application to MLLMs faces fundamental challenges. The inherent heterogeneity of multimodal data introduces distinct redundancy patterns absent in unimodal text. These include, but are not limited to, spatial correlations in high-resolution images, spatiotemporal continuity in video sequences, and spectral-temporal locality in audio streams. Such specialized redundancies necessitate the development of dedicated compression strategies. Consequently, this survey systematically reviews emerging token compression methodologies for MLLMs that effectively reduce token redundancy while preserving task performance.
2.3 Vision Transformer Token Compression
@@
Visual token compression, originally pioneered in vision transformers (ViTs) (Vaswani et al., 2017; Dosovitskiy et al., 2020; Dong et al., 2022; Liu et al., 2021; Fan et al., 2021; Li et al., 2022; Graham et al., 2021; Huang et al., 2025a; Feng and Zhang, 2023), offers insights for addressing analogous challenges in MLLMs.
@@
Spatial redundancy manifests in ViTs through adjacent image patches, where not all tokens contribute equally to classification outcomes, compounded by semantic imbalance: foreground objects demand disproportionate computational resources compared to homogeneous backgrounds. To mitigate these issues, visual token compression techniques are employed to reduce computational overhead while maintaining model accuracy.
@@
Foundational approaches, including DynamicViT (Rao et al., 2021) and EViT (Liang et al., 2022), quantify token relevance through attention scores, dynamically pruning low-saliency tokens. Complementary techniques like ToMe (Bolya et al., 2022) and TokenLearner (Ryoo et al., 2021) either merge semantically similar tokens using similarity metrics or generate compact token sets via learned spatial attention mechanisms. DeiT (Touvron et al., 2021) employs lightweight ‘student’ heads to predict categorical labels from compressed token subsets. Furthermore, methods such as MADTP (Cao et al., 2024) leverage cross-modal alignment to filter tokens.
@@
2.4 Problem Definition and Taxonomy Scope
To clarify the scope of this survey and distinguish token compression from related efficient computing techniques, we establish a strict criterion based on the physical reduction of information flow. We define a method as token compression if and only if it explicitly reduces the number of tokens passed to subsequent layers or modules.
Formally, given an input sequence $\mathbf{X}\in\mathbb{R}^{N\times D}$, a token compression operator $\mathcal{T}$ produces an output $\mathbf{X}^{\prime}\in\mathbb{R}^{M\times D}$ where $M<N$, while aiming to retain the essential semantic information of $\mathbf{X}$. Based on this definition, we delineate the boundaries with related concepts as follows:
  • Input-level Compression: We classify techniques such as frame sampling and key-frame extraction as a generalized form of token compression operating at the Input Level. By selecting a subset of frames (e.g., extracting key-frames from a video), the initial token count $N$ is reduced prior to the encoding stage. We distinguish this from Feature-level Compression (e.g., token pruning or merging), which dynamically operates on intermediate embeddings within the network layers.
  • Exclusion of Attention Sparsity: We exclude attention sparsity mechanisms (and related efficient attention variants) from the scope of token compression. While these methods reduce computational complexity (e.g., from $O(N^{2})$ to linear) by masking interactions, they typically output a sequence of the same length ($N\to N$) to the next layer. They sparsify the computation graph, whereas token compression sparsifies the representation.
Table 1: Four Categories of Methods Based on Intrinsic Mechanisms: Diagram, Summary, and Pros & Cons.